40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 21 11:55:05 2015, Steve, Update, Optical Levers, 1103P noise measurement 1103P@qpdBS.jpgoplev1103P.pngoplevSum1103Pbs.png
    Reply  Tue Apr 21 15:03:27 2015, rana, Update, Optical Levers, 1103P noise measurement 
    Reply  Tue Apr 21 21:46:34 2015, rana, Update, Optical Levers, 1103P noise measurement oplev1103P.pngL1-ISI_SUS_ODC_FC496F_SPECTRUM-1113523216-86400.png
       Reply  Wed Apr 22 11:21:51 2015, Steve, Update, Optical Levers, BS & PRM oplevs are back to normal 
Message ID: 11230     Entry time: Tue Apr 21 11:55:05 2015     Reply to this: 11231   11232
Author: Steve 
Type: Update 
Category: Optical Levers 
Subject: 1103P noise measurement 

Manasa and Steve,

Is this what you want?  Dashed lines are dark.

BS and PRM oplevs are blocked for this measurement. I will restore to normal operation at 4pm today.
 

Attachment 1: 1103P@qpdBS.jpg  357 kB  Uploaded Tue Apr 21 15:13:02 2015  | Hide | Hide all
1103P@qpdBS.jpg
Attachment 2: oplev1103P.png  87 kB  Uploaded Tue Apr 21 15:14:48 2015  | Hide | Hide all
oplev1103P.png
Attachment 3: oplevSum1103Pbs.png  76 kB  Uploaded Tue Apr 21 15:15:20 2015  | Hide | Hide all
oplevSum1103Pbs.png
ELOG V3.1.3-